# GPTVoiceTasker: LLM-Powered Virtual Assistant for Smartphone

### v1

@@ Abstract.
Virtual assistants have the potential to play an important role in helping users achieves different tasks. However, these systems face challenges in their real-world usability, characterized by inefficiency and struggles in grasping user intentions. Leveraging recent advances in Large Language Models (LLMs), we introduce GptVoiceTasker, a virtual assistant poised to enhance user experiences and task efficiency on mobile devices. GptVoiceTasker excels at intelligently deciphering user commands and executing relevant device interactions to streamline task completion. The system continually learns from historical user commands to automate subsequent usages, further enhancing execution efficiency. Our experiments affirm GptVoiceTasker’s exceptional command interpretation abilities and the precision of its task automation module. In our user study, GptVoiceTasker boosted task efficiency in real-world scenarios by 34.85%, accompanied by positive participant feedback. We made GptVoiceTasker open-source, inviting further research into LLMs utilization for diverse tasks through prompt engineering and leveraging user usage data to improve efficiency.
@@
copyright: noneccs: Human-centered computing Interaction techniquesccs: Human-centered computing Smartphonesccs: Human-centered computing Natural language interfacesccs: Human-centered computing Sound-based input / output
Figure 1. An example use case in Home Workout application when the user needs to interact with the smartphone hands-free due to physical busyness. When performing an on-screen interaction, GptVoiceTasker repeatedly prompts LLM to predict on-screen actions with current UI information and execute the response to achieve user tasks. The execution information is then saved to streamline the execution of subsequent similar tasks.
Figure 1. An example use case in Home Workout application when the user needs to interact with the smartphone hands-free due to physical busyness. When performing an on-screen interaction, GptVoiceTasker repeatedly prompts LLM to predict on-screen actions with current UI information and execute the response to achieve user tasks. The execution information is then saved to streamline the execution of subsequent similar tasks.
@@ 1. Introduction
The advancements in voice control technology have sparked a new wave of innovation, driving the exploration of its potential in transforming smartphone interactions (Hoy, 2018; Hirschberg and Manning, 2015). With the integration of voice control, users can effortlessly navigate through various applications, compose messages, and even initiate tasks like checking the weather or playing a video on YouTube (Vu et al., 2023). This seamless and natural mode of interaction not only saves time but also promotes a hands-free experience, allowing individuals to engage with their smartphones in situations where manual input operation is impractical or inconvenient (Zhong et al., 2014; Liu et al., 2015) (see Fig. 1). Moreover, the success stories of widely recognized voice assistants like Google Voice Assistant (Assistant, 2022) and Siri (Apple, [n. d.]) have further propelled the adoption of this technology, inspiring researchers and developers to delve deeper into its capabilities and refine its usability for an even broader range of users. As this field continues to evolve, the integration of voice control technology with smartphones holds immense potential to revolutionize the way we interact with our devices, empowering users with enhanced accessibility and a more convenient user experience.
Developing efficient and reliable voice-controlled systems involves addressing various challenges that significantly impact the accuracy and usability of these systems (Myers et al., 2018). One major hurdle is accurately comprehending user commands and seamlessly mapping them to specific smartphone actions (Pan et al., 2022). Achieving satisfactory accuracy levels often relies on substantial training data, comprised of user voice commands, and complex deep learning models (Li et al., 2020a). These models are employed to decipher the semantic nuances of the commands, comprehend their intended meaning, and subsequently select the appropriate user interface (UI) element on the screen for interaction. However, acquiring and curating substantial datasets for training language models presents a laborious and time-consuming challenge, impeding progress and scalability in developing virtual assistants. Moreover, the collected dataset may not comprehensively cover the diverse commands required for modern applications with rich functionalities, potentially limiting the language model’s capabilities. Additionally, the rapid advancement of technology constantly introduces new functionalities, quickly rendering the existing dataset outdated, and these models often struggle with accommodating the interpreting speech recognizer errors due to accents (Tadimeti et al., 2021) and homophones (Nadkarni et al., 2011). The inherent complexity of these models hampers their adaptability to the dynamic landscape of human natural language and pronunciation variations, making it challenging to account for real-world usage scenarios’ intricacies and variations.
Furthermore, voice assistants often face performance issues when compared to physical touches on mobile devices (Angelini et al., 2016). Extensive research and development have explored various methods to enhance voice assistants’ capabilities. For instance, Voicify (Vu et al., 2023) employs app analysis and deep linking techniques to speed up voice command processing. Modern voice assistants also enhance user experience by personalizing interactions based on individual usage patterns and preferences (Braun and Alt, 2019; Yuan et al., 2019). This personalized learning approach holds tremendous potential in mitigating the performance limitations of voice-controlled systems, enhancing efficiency, and accelerating user interactions (Li et al., 2017; Pan et al., 2022). By continuously learning from user interactions, voice assistants can improve their accuracy and reliability, ultimately ensuring efficiency and overall user experience in real-world usage scenarios.
Large Language Models (LLMs) have revolutionized natural language processing (NLP), showcasing their general intelligence and excellence in reading comprehension, trivia quizzes, translation, and text completion (Bareiß et al., 2022). The concept of Few-Shot Learning further enhances LLMs’ power by enabling them to adapt to new tasks with minimal examples or prompts, eliminating the need for task-specific models and datasets (Perez et al., 2021). This flexibility and efficiency in handling diverse conversational interactions without extensive retraining present an innovative and promising approach. Leveraging few-shot learning, developers can tap into the vast knowledge and language understanding capabilities of LLMs, facilitating rapid development and deployment of language-based automation solutions in various domains (Liu et al., 2023b; Fan et al., 2023). Therefore, we aim to investigate the potential of LLMs in enhancing the intuitiveness of voice control for mobile applications.
This paper introduces GptVoiceTasker, a novel system that leverages LLMs to enhance voice control for Android devices. Our system employs advanced prompt engineering techniques to ensure a precise understanding of user commands without extensive model training. By bridging the gap between natural language commands and interactive tasks on mobile devices, GptVoiceTasker facilitates seamless automation of daily physical interactions, including scrolling, tapping, and inputting text, solely through voice commands. In addition, the system automatically records and learns from user commands and in-app usages, enabling the reproduction of tasks for similar requests in the future to improve task efficiency. GptVoiceTasker supports the sequential execution of multiple commands within a single utterance, further enhancing the efficiency of voice interaction on mobile devices.
@@
We validated the technical contributions of GptVoiceTasker by evaluating i) the ability to parse user commands into executable actions and ii) the ability to streamline saved tasks. The command parser achieved over 90% accuracy on the human command dataset collected from a user study. Our automated execution achieved 82.7% success rate for direct match tasks and 72.0% success rate for tasks with different parameters. To validate the usability of GptVoiceTasker, we conducted a user evaluation with 18 participants, each completing a set of tasks using GptVoiceTasker and two state-of-the-art baselines. We collected the time taken to complete each task, as well as quantitative and quality feedback from users. The results show that GptVoiceTasker accelerates the tasks by 34.85% and received positive feedback regarding usability.
@@
  • Conducting a large-scale user evaluation to validate the effectiveness of our approach, along with empirical findings on system limitations and considerations for voice assistant design.
@@ 2. Background & Related works
2.1. Voice Control & Assistants on Mobile Devices
@@
In recent years, the progress in Natural Language Understanding (NLU) has fueled the proliferation of voice assistants across diverse platforms, including ubiquitous systems (Krishna and Nagendram, 2012; Bai, 2022) and home appliances (Park and Kim, 2018). On personal smartphones, numerous voice assistants have been proposed, leveraging intelligent NLU models to comprehend user speech input and map it to user interface actions. An early milestone in smartphone voice control interfaces was JustSpeak, which harnessed Google’s Automatic Speech Recognition (ASR) to record user commands and introduced innovative utterance parsing techniques (Zhong et al., 2014). Subsequently, the Smart Voice Assistant expanded on JustSpeak’s capabilities by enabling users to manage calls and SMS through voice commands (Bhalerao et al., 2017). A similar endeavor, the "VoiceNavigator" application by Weber et al. in 2016, focused on enhancing the visibility and learnability of mobile voice user interface applications (Corbett and Weber, 2016). However, these initial approaches, foundational as they were, encountered usability issues stemming from rigid language parsing heuristics and limited use cases, which spurred the need for further development of smartphone virtual assistants.
@@
In recent years, significant advancements in language parsing capabilities have been achieved through deep learning models. SAVANT leveraged Dialogflow as a conversational agent to extract user intent from utterances (Arsan et al., 2021), while DoThisHere employed the pre-built Almond language model to enable voice control for retrieving and setting UI contents in Android (Yang et al., 2020). Google released Voice Access (Access, 2022), aimed to replace manual interactions with voice command, which has over 100 millions downloads on Google Play Store. More recently, Voicify (Vu et al., 2023) introduced VoicifyParser, an advanced deep learning approach for parsing user commands into on-screen interactions, while AutoVCI (Pan et al., 2022) concentrated on developing voice interfaces for automating mobile UI tasks. However, the interaction paradigm with these existing approaches remains somewhat unnatural, requiring users to issue precise machine-like instructions, such as “Press save button”. This limitation means that they may struggle to fully comprehend high-level user intentions, such as “I want to save this note”. We propose GptVoiceTasker to address these challenges and revolutionise the voice-based interactions between human and software systems. Our solution leverages the capabilities of LLMs to map high-level user intentions to executable actions, enabling on-screen interactions through intention-based voice commands. This approach seeks to accommodate the flexibility and natural language of human commands, ushering in a new era of user-friendly assistive tools.
2.2. Large Language Models for Enhanced Human-AI Collaboration
@@
The advent of generative AI has given rise to innovative LLMs, such as GPT-4 (OpenAI, 2023b) and DALL-E (Ramesh et al., 2021). These LLMs have revolutionized the landscape of AI development by enabling developers to achieve complex tasks through few-shot prompting, eliminating the need for extensive custom model training. Their remarkable versatility has spurred active research in both IT and non-IT domains, spanning areas like software testing (Liu et al., 2023a; Feng and Chen, 2023), high-performance computing (Chen et al., 2023), finance (Wu et al., 2023), and health science (Gilbert et al., 2023). LLMs have particularly excelled in enhancing the intuitiveness of existing methods, as seen in software testing, where they generate authentic text inputs based on the current UI page information, replacing the conventional random text input approach (Liu et al., 2023a). This demonstrates the transformative potential of LLMs in advancing research and innovation across a multitude of domains.
@@
The capabilities of LLMs have sparked a surge in their application within assistive technology, revolutionizing the translation of user commands into executable tasks across diverse systems. Recent research in this domain has witnessed the transformation of human natural language commands into various types of tasks, including visualization tasks (Wang et al., 2023b), operating system tasks (Liu et al., 2023e), and robotic tasks (Singh et al., 2023; Liu et al., 2023c). LLMs have enabled these systems to tackle more intricate commands beyond the scope of existing heuristic approaches. They also exhibit a remarkable ability to comprehend variations of commands that share similar intentions but are expressed differently. This pioneering framework, with the support of LLMs, paves the way for a novel (semi)automated task execution paradigm, erasing the boundaries between traditional command patterns and intuitive command modalities.
Within the domain of mobile assistants, LLMs have emerged as a transformative force, supplanting conventional machine learning models as seen in prior related works (Vu et al., 2023; Pan et al., 2022). This paradigm shift simplifies the process of translating user’s natural voice commands into actionable operations on mobile User Interfaces (UIs). Wang et al. (Wang et al., 2023a) utilised LLMs to allow conversation-alike interaction with mobile UIs, which demonstrates LLM’s better ability in understanding on-screen elements compared with traditional machine learning approaches (Li et al., 2020b). While Wang et al.’s approach in smartphone virtual assistant development is noteworthy, its employed prompting strategy remains basic, impacting its ability to parse natural user commands. Our work aims to elevate this approach by incorporating state-of-the-art prompting strategies to enhance accuracy and create a more comprehensive smartphone virtual assistant. Unlike Wang et al.’s focus on single on-screen actions, our research takes a step further. We strive to unlock the full potential of LLMs in automating complex user requests involving multiple steps to achieve specific objectives, as exemplified in Fig. 1 with the query "Show me how to do the chest fly exercise". Our goal is to broaden the scope of LLMs, assisting users not only in known tasks but also in accomplishing unfamiliar ones, all while significantly reducing the time required to complete these tasks on a smartphone.

### v2

# GPTVoiceTasker: Advancing Multi-step Mobile Task Efficiency Through Dynamic Interface Exploration and Learning

@@ Abstract.
Virtual assistants have the potential to play an important role in helping users achieves different tasks. However, these systems face challenges in their real-world usability, characterized by inefficiency and struggles in grasping user intentions. Leveraging recent advances in Large Language Models (LLMs), we introduce GptVoiceTasker, a virtual assistant poised to enhance user experiences and task efficiency on mobile devices. GptVoiceTasker excels at intelligently deciphering user commands and executing relevant device interactions to streamline task completion. For unprecedented tasks, GptVoiceTasker utilises the contextual information and on-screen content to continuously explore and execute the tasks. In addition, the system continually learns from historical user commands to automate subsequent task invocations, further enhancing execution efficiency. From our experiments, GptVoiceTasker achieved 84.5% accuracy in parsing human commands into executable actions and 85.7% accuracy in automating multi-step tasks. In our user study, GptVoiceTasker boosted task efficiency in real-world scenarios by 34.85%, accompanied by positive participant feedback. We made GptVoiceTasker open-source, inviting further research into LLMs utilization for diverse tasks through prompt engineering and leveraging user usage data to improve efficiency.
@@
copyright: acmlicensedconference: ACM Symposium on User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA, USAdoi: 10.1145/3654777.3676356ccs: Human-centered computing Interaction techniquesccs: Human-centered computing Smartphonesccs: Human-centered computing Natural language interfacesccs: Human-centered computing Sound-based input / output
Figure 1. GptVoiceTasker provides an intuitive method for automating complex commands on smartphones during physically demanding activities, such as cooking. It automatically explores step-by-step interactions to complete unprecedented tasks and uses the saved information to accelerate the automation process for tasks that have been previously encountered.
@@ 1. Introduction
The advancements in voice control technology have sparked a new wave of innovation, driving the exploration of its potential in transforming smartphone interactions (Hoy, 2018; Hirschberg and Manning, 2015). With the integration of voice control, users can effortlessly navigate through various applications, compose messages, and even initiate tasks like checking the weather or playing a video on YouTube (Vu et al., 2023). This natural mode of interaction not only saves time but also promotes a hands-free experience, allowing individuals to engage with their smartphones in situations where manual input operations are impractical or inconvenient (Zhong et al., 2014; Liu et al., 2015) (refer to Figure. 2 illustrating a user engaged in gym activities). Moreover, the success stories of widely recognized voice assistants like Google Voice Assistant (Assistant, 2022) and Siri (Apple, [n. d.]) have further propelled the adoption of voice-based interactions, inspiring researchers and developers to delve deeper into its capabilities and refine its usability for an even broader range of users.
Developing autonomous voice-controlled assistants involves addressing various challenges that impact the usability of these systems (Myers et al., 2018). Current industrial products (such as Siri and Google Assistant) do not provide a universal approach as they require app developers to explicitly define a narrow set of voice-supported actions within the code. For instance, while those voice assistants fully support searching for videos on YouTube, other YouTube features such as accessing video history is unavailable. This results in a disjointed and often unfinished experience for users as they cannot rely on voice interactions to fully control their smartphone. In addition, these assistants often struggle to comprehend commands when users’ inputs are unclear or do not align with the predefined patterns (Developers, 2024b, a). This lack of understanding further impacts the overall utility of the voice-controlled systems. Furthermore, app developers face the challenge of anticipating and programming an extensive variety of potential intents, a task that is both complex and limiting in the scope of voice assistant capabilities.
Recently, Large Language Models (LLMs) have brought a paradigm shift in natural language processing (NLP), demonstrating remarkable capabilities in tasks like reading comprehension, translation, and text completion (Bareiß et al., 2022). The advent of Few-Shot Learning has further amplified the capabilities of LLMs, enabling them to quickly adapt to new logical reasoning tasks with minimal examples. This versatility and efficiency in managing various conversational interactions without the need for extensive retraining offer a groundbreaking approach, obviating the need for task-specific models and extensive datasets (Perez et al., 2021). In the context of mobile assistants/agents, the integration of LLMs to comprehend user commands and interface with mobile UIs has gained traction (Wen et al., 2023; Wang et al., 2023a; Yan et al., 2023; Pan et al., 2023; Carreira et al., 2023). Amazon’s integration of LLMs into Alexa represents a significant step in enhancing voice assistants (Developer, 2024). Their primary focus has been on improving Alexa’s ability to understand users’ needs more accurately and to control other devices more effectively. But it’s not clear how Alexa can help with smartphones users and if the assistant is optimized with mobile UIs. Some research in this area concentrates on translating mobile GUIs into text, relying on LLMs to understand the context and predict interactive screen elements (Wang et al., 2023a; Wen et al., 2023). However, this method sometimes struggles with the extraction of irrelevant GUI elements or fails when the command does not directly relate to the current screen. Other UI automation tools for UI testing explored enhancing LLMs with image processing capabilities, such as those found in GPT-4v (Yan et al., 2023). While this approach shows a higher success rate, it is hampered by longer processing times and increased costs, which can negatively impact user experience in real-time systems like voice assistants.
This paper introduces GptVoiceTasker, a novel voice assistant that automates multi-step unprecedented tasks by dynamically exploring app interface and accelerate similar tasks through prior usages. Drawing inspiration from the conventional record-and-replay approach (Lam et al., 2017; Li et al., 2017), GptVoiceTasker is designed to learn GUI transformations as users navigate through apps. Instead of simply transmitting live UI information, GptVoiceTasker captures information of the current UI as the user interacts with the app, storing this data in a backend database as known knowledge to the system. Whenever a user command is issued, GptVoiceTasker cross-references the new executing task with this stored knowledge to make informed decisions. This method allows reproduction of tasks for similar future requests, enhancing task efficiency and accuracy. If a user command references a unprecedented feature, GptVoiceTasker will step-by-step explore and record the navigational path to achieve the feature. Our system achieves this by advanced prompt engineering techniques to ensure a precise understanding of user commands without extensive model training. GptVoiceTasker effectively bridges the divide between natural language commands and interactive mobile tasks, enabling seamless automation of everyday tasks that include actions like scrolling, tapping, and text input purely via voice commands.
@@
We validated the technical contributions of GptVoiceTasker by evaluating i) the ability to parse user commands into executable actions, ii) the ability to complete multi-step tasks given one command, and iii) the ability to streamline saved tasks. The command parser achieved over 90% accuracy on a human command dataset collected from the user study. GptVoiceTasker also achieved 85.7% success rate in completing human-collected multi-step tasks. Our automated execution achieved 82.7% success rate for direct match tasks and 72.0% success rate for tasks with different parameters. To validate the usability of GptVoiceTasker, we conducted a user evaluation with 18 participants, each completing a set of tasks using GptVoiceTasker and two state-of-the-art baselines. We collected the time taken to complete each task, as well as quantitative and quality feedback from users. The results showed that GptVoiceTasker accelerated the tasks by 34.85% and received positive feedback regarding usability.
@@
  • Conducting a user evaluation to validate the effectiveness of our approach, along with empirical findings on system limitations and considerations for voice assistant design.
@@ 2. Background & Related works
2.1. Voice Control & Automation on Mobile Devices
Recent advancements in Natural Language Understanding (NLU) have significantly enhanced the development of voice assistants across various platforms, including ubiquitous systems (Krishna and Nagendram, 2012; Bai, 2022) and home appliances (Park and Kim, 2018).
@@ 2.1. Voice Control & Automation on Mobile Devices
An early milestone in smartphone voice control interfaces was JustSpeak, which harnessed Google’s Automatic Speech Recognition (ASR) to record user commands and introduced innovative utterance parsing techniques (Zhong et al., 2014). Subsequently, the Smart Voice Assistant expanded on JustSpeak’s capabilities by enabling users to manage calls and SMS through voice commands (Bhalerao et al., 2017). However, these initial approaches, foundational as they were, encountered usability issues stemming from rigid language parsing heuristics and limited use cases, which spurred the need for further development of smartphone virtual assistants.
@@
In recent years, significant advancements in language parsing capabilities have been achieved through deep learning models. SAVANT leveraged Dialogflow as a conversational agent to extract user intent from utterances (Arsan et al., 2021), while DoThisHere employed the pre-built Almond language model to enable voice control for retrieving and setting UI contents in Android (Yang et al., 2020). Google released Voice Access (Access, 2022), aimed to replace manual interactions with voice command, which has over 100 millions downloads on Google Play Store. Moreover, there are Firefox Voice, an open and extensible web-based voice assistant with speech-to-text engine (Cambre et al., 2021), and Talk2Care, which leverages LLMs to facilitate communication between healthcare providers and older adults (Yang et al., 2024). Additionally, Just Speak It has focused on minimizing cognitive load during eyes-free text editing with a smart voice assistant (Fan et al., 2021), while GazePointAR uses context-aware multimodal inputs for pronoun disambiguation in augmented reality (Lee et al., 2024). Voicify (Vu et al., 2023) introduced VoicifyParser, an advanced deep learning approach for parsing user commands into on-screen interactions. However, the interaction paradigm with these existing approaches remains somewhat unnatural, requiring users to issue precise machine-like instructions, such as “Press save button”. This limitation means that they may struggle to fully comprehend high-level user intentions, such as “I want to save this note”. We propose GptVoiceTasker to address these challenges and revolutionise the voice-based interactions between human and software systems. Our solution leverages the capabilities of LLMs to map high-level user intentions to executable actions, enabling on-screen interactions through intention-based voice commands. This approach seeks to accommodate the flexibility and natural language of human commands, ushering in a new era of user-friendly assistive tools.
Research has also delved into voice command interfaces for automating smartphone tasks, often categorized as programming-by-demonstration tools (Arsan et al., 2021). These systems generally utilize a record-and-replay strategy, where the user records a series of actions to complete a task and later triggers that sequence with a voice command. SUGILITE (Li et al., 2017), for instance, introduced methods for performing task variations with different parameters from a single recorded instance. Building upon this, AutoVCI (Pan et al., 2022) automated the generation of verbal commands for activating saved tasks. However, these tools entail usability challenges as they require users to manually record execution paths for each use case. In contrast, GptVoiceTasker innovates by predicting the most suitable action for each UI screen based on user requests without the need for pre-programming. Our database, tailored for streamlining task execution, is automatically constructed in real-time as users interact with their mobile apps.
$2.$ Large Language Models for Enhanced Human-AI Collaboration
@@
The advent of generative AI has given rise to innovative LLMs, such as GPT-4 (OpenAI, 2023b), DALL-E (Ramesh et al., 2021) and Llama (Touvron et al., 2023). These LLMs have revolutionized the landscape of AI development by enabling developers to achieve complex tasks through few-shot prompting, eliminating the need for extensive custom model training. Their remarkable versatility has spurred active research in both IT and non-IT domains, spanning areas like software testing (Liu et al., 2023a; Feng and Chen, 2023), high-performance computing (Chen et al., 2023), finance (Wu et al., 2023), and health science (Gilbert et al., 2023). LLMs have particularly excelled in enhancing the intuitiveness of existing methods, as seen in software testing, where they generate authentic text inputs based on the current UI page information, replacing the conventional random text input approach (Liu et al., 2023a). This demonstrates the transformative potential of LLMs in advancing research and innovation across a multitude of domains.
@@
The capabilities of LLMs have sparked a surge in their application within assistive technology, revolutionizing the translation of user commands into executable tasks across diverse systems. Recent research in this domain has witnessed the transformation of human natural language commands into various types of tasks, including visualization tasks (Wang et al., 2023b), operating system tasks (Liu et al., 2023d), and robotic tasks (Singh et al., 2023; Liu et al., 2023c). LLMs have enabled these systems to tackle more intricate commands beyond the scope of existing heuristic approaches. They also exhibit a remarkable ability to comprehend variations of commands that share similar intentions but are expressed differently. This pioneering framework, with the support of LLMs, paves the way for a novel (semi)automated task execution paradigm, erasing the boundaries between traditional command patterns and intuitive command modalities.
Notable advancements include the World of Bits platform, which enables web-based agent training through interactions with real-world websites using low-level actions. WoB utilizes reinforcement learning and behavioral cloning to demonstrate these techniques’ potential in web-based tasks, ensuring reproducibility through cached HTTP traffic (Shi et al., 2017). Additionally, Mind2Web pushes the boundaries of generalist web agents by leveraging LLMs to handle complex, open-ended tasks across a wide range of real-world websites, showcasing the capacity of LLMs to generalize across diverse web environments (Deng et al., 2024). Generative agents further highlight the application of LLMs in simulating human behavior in interactive settings, providing a framework for more dynamic and lifelike simulations (Park et al., 2023). Lastly, the design framework involving cells, generators, and lenses aims to optimize object-oriented interactions with LLMs, enhancing usability and functionality in various applications (Kim et al., 2023).
Within the domain of mobile assistants, LLMs have become a transformative force, overtaking traditional machine learning models as demonstrated in previous works (Vu et al., 2023; Pan et al., 2022). This shift has simplified the translation of natural voice commands into actions on mobile UIs. Wang et al. (Wang et al., 2023a) used LLMs for conversation-like interactions with mobile UIs, showcasing a superior understanding of on-screen elements compared to earlier machine learning methods (Li et al., 2020). However, their approach only focuses on single-screen support and interactions, which is inadequate for completing multi-step tasks. AutoDroid (Wen et al., 2023) and AutoTask (Pan et al., 2023) have employed LLMs, incorporating a degree of application knowledge (e.g., recalling previous actions or repeating similar commands) to execute multi-step tasks through a single command. that can complete multi-steps tasks under one command. Yet, these methods have tended to concentrate on discrete tasks without fully addressing the continuity between tasks within the same application. GptVoiceTasker advances this field by collecting sophisticated domain knowledge and employing advanced prompt techniques, aimed at improving precision and establishing a more advanced smartphone virtual assistant. This enhancement allows users to execute both familiar and novel tasks more effectively on their devices.